# Gradient Minimization Using Higher Order MethodsYao Ji and Guanghui Lan

### v1

@@ 3.1 Extension from accelerated cubic newton to high-order tensor method
For the $S$-th epoch, there exists $v_{S}^{k}\in\partial f_{S}({x}_{S}^{k}),N_{S}\leq k\leq 2N_{S},$ such that
@@
Furthermore, in the last epoch $S,$ for all $N_{S}\leq k\leq 2N_{S},$ there holds $\|x_{S}^{k}-x_{S}^{*}\|\leq{\|x_{S-1}-x_{S-1}^{*}\|}/{4}.$
@@
Combining it with (19) in 1 and (17) in Lemma 3.2, we have
@@
Similarly, notice that in the last epoch, for all $N_{S}\leq k\leq 2N_{S},$ there holds
@@
Denote $p$-th order oracle at point $x$ as $\{f(x),\nabla f(x),\nabla^{2}f(x),\dots,\nabla^{p}f(x)\},p\geq 1.$ The following theorem states the main convergence results of the high-order AR framework with subroutine $\mathcal{A}$ applied for convex composite problem (3).
@@
(26) $\tfrac{16\left[2(p+\nu){C}_{\mathcal{A}}L_{p,\nu}(g)D^{p+\nu-1}\right]^{\frac{1}{p+\nu}}}{\left(1-2^{-1+\frac{1}{p+\nu}}\right)\varepsilon^{\frac{1}{p+\nu}}}=\Theta\left(\tfrac{L_{p,\nu}(g)^{\frac{1}{p+\nu}}D^{\frac{p+\nu-1}{p+\nu}}}{\varepsilon^{\frac{1}{p+\nu}}}\right)$
calls to the $p$-th order oracle.
@@
In view of Theorem 3.6, we can accelerate the convergence rate of the gradient norm by employing various tensor-based methods as subroutines. Examples include the accelerated cubic Newton method [23], accelerated tensor methods [25], the bi-level unconstrained minimization approach [28], and accelerated proximal-point methods with lower-order solvers [27, 28]. Notably, in [27, 28], they show that second-order methods can achieve the same convergence rate as third-order methods in terms of the function value residual. As a result, using such second-order methods as subroutines in the high-order AR framework Algorithm 3 allows the convergence rate in terms of the gradient norm to match that of the function value residual in [27, 28]. In what follows, we present the Accelerated Regularized Tensor Method (ARTM) [15, 14] as an inner subroutine for solving (3) with $\nu\in[0\,1]$, which improves the gradient complexity bound of ARTM in [15, Theorem 5.7] by removing a logarithmic factor. The gradient complexities for the other methods follow by the same reasoning.
@@
Suppose the subroutine $\mathcal{A}$ is ARTM. Then, Algorithm 3 can compute an approximate solution $x_{S}^{k}$ s.t., $\min_{k=N_{S}+1,\dots,2N_{S}}\|\nu_{S}^{k}\|\leq\varepsilon$, where $\nu_{S}^{k}\in\nabla g(x_{S}^{k})+\partial h(x_{S}^{k})$, within $\mathcal{O}(({D^{{p+\nu-1}{}}}/{\varepsilon})^{\frac{1}{p+\nu}})$ calls to the $p$-th order oracle, where $D$ is defined in (16), and $p\geq 2,$ $\nu\in[0\,\,1].$
@@
[15, Theorem 5.7] shows that 1 holds for ARTM, thus Theorem 3.6 applies.
This result is the tightest known for convex and $p$-times differentiable fucntion with $\nu$-Hölder continuous $p$-th derivative. Although it does not match the lower complexity bounds for such function class  [15, Theorem 6.6], given by $\Omega\left(\left({D^{{p+\nu-1}}}/{\varepsilon}\right)^{\frac{2}{3(p+\nu)-2}}\right).$ The gap arises because no existing methods have yet achieved the optimal rate in terms of the function value residual for such function class. If such methods are developed, the high-order AR framework can directly incorporate them as subroutines $\mathcal{A}$ to accelerate the gradient norm convergence and achieve a matching rate.
@@ 3.2 Fast gradient minimization for convex smooth function
In this subsection, we study AR with optimal $p$-th order methods as subroutines to handle composite convex functions with Lipschitz continuous $p$-th derivatives. Specifically, there exists $L_{p+1}(g)>0$ s.t.,
@@
For the $S$-th epoch, there exists $v_{S}^{k}\in\partial f_{S}({x}_{S}^{k}),N_{S}\leq k\leq 2N_{S},$ such that
@@
Compared to 1, 2 imposes a faster convergence rate in terms of both the function gap and the gradient norm. Observe that the gradient norm convergence rate still is slower than the function value residual. The above Assumption holds for a wide range of optimal algorithms with different $p.$ The subroutines examples include first order algorithm: the accelerated gradient descent method (AGD) [29], second order algorithm: large stepsize accelerated hybrid proximal extragradient [22], and $p$-th order algorithm: optimal tensor method (OTM) [6, 19, 12] or inexact $p$-th order proximal point method with tensor step [26, 28].
@@
where $D$ is defined in (16). If the subroutine for solving subproblems satisfies 2. Then, for all $s\in[S],$ Algorithm 3 can compute $x_{s}$ such that $\|x_{s}-x_{s}^{*}\|\leq{\|x_{s-1}-x_{s-1}^{*}\|}/{4}.$ Furthermore, in the last epoch $S,$ for all $N_{S}\leq k\leq 2N_{S},$ there holds $\|x_{S}^{k}-x_{S}^{*}\|\leq{\|x_{S-1}-x_{S-1}^{*}\|}/{4}.$
@@
(32) $\tfrac{8\left[(p+1)L_{p+1}(g)D^{p}\right]^{\frac{2}{3p+1}}}{\varepsilon^{\frac{2}{3p+1}}}=\Theta\left(\left[\tfrac{L_{p+1}(g)D^{p}}{\varepsilon}\right]^{\frac{2}{3p+1}}\right)$
calls to the $p$-th order oracle.
@@
Suppose the subroutine $\mathcal{A}$ is A-NPE [22]. Then, a third-order AR framework Algorithm 3 can compute an approximate solution $x_{S}^{k}$ such that $\min_{k=N_{S}+1,\dots,2N_{S}}\|\nu_{S}^{k}\|\leq\varepsilon$, where $\nu_{S}^{k}\in\nabla g(x_{S}^{k})+\partial h(x_{S}^{k})$, within
@@
[22, Theorem 4.1] shows that 2 holds for A-NPE, thus Theorem 3.11 applies.
We next show a gradient norm minimization using a second-order method for a third-order smooth problem, and show its convergence rate of the gradient norm matches that of the function value residual in [26]. For simplicity, we consider the unconstrained problem (1) in this example.
@@
Suppose the subroutine $\mathcal{A}$ is the second order implementation of the inexact third-order proximal point method [26]. Then, a fourth-order AR Algorithm 3 can compute an approximate solution $x_{S}^{k}$ such that
@@
within $\mathcal{O}([{L_{4}(f)D^{3}}/{\varepsilon}]^{\frac{1}{5}})$ iterations, where $D$ is defined in (16), at each iteration, it calls the second-order and the first-order oracle $\mathcal{O}(\log{\varepsilon^{-1}})$ times at most.
@@
[26, Theorem 3.8, 5.5] show that 2 holds with $p=3$, thus Theorem 3.11 applies.
@@ 4 Parameter-free algorithm for convex problems
For all convergence results in the Section 2 and Section 3 are based on the exact solutions of the auxiliary optimization problem at each iteration of the high-order method, and assumed to have access to certain key problem parameters, including the Lipschitz constant of the $\nu$-Hölder continuous $p$-th derivative $L_{p,\nu}(g)$, and an upper bound $D$ on the initial distance to the solution set, such that $\text{dist}(x_{0},X^{*})\leq D$. However, in generally, auxiliary optimization problem at each iteration can not be solved in general, and those problem dependent parameters are not known before hand. In this section, we investigate inexact and parameter-free implementations of high-order methods for gradient minimization that do not rely on such information. For clarity of exposition, we first focus on the unconstrained setting; analogous parameter-free strategies can be developed for problems with simple constraints composite structures.
Algorithm 4 High-order AR without the knowledge of $L_{p,\nu}$
1: Initial regularization $\sigma_{0}=0,\sigma_{1}>0,$ and initial point $x_{0}\in\mathbb{R}^{n},$ and $L_{0}.$
2: for $s=1,2\dots$ do
@@
4:   Compute an approximate solution $x_{s}$ of the proximal subproblem (33) $x_{s}\approx\operatorname*{arg\,min\,}\limits_{x\in\mathbb{R}^{n}}\left\{f_{s}(x)\coloneqq f(x)+\textstyle\sum_{i=1}^{s}\tfrac{\sigma_{i}-\sigma_{i-1}}{p+\nu}\|x-{x}_{i-1}\|^{p+\nu}\right\},$
5:   where $p\geq 1,\nu\in[0\,\,1],$ and $p+\nu\geq 2$ by running $\mathcal{A}_{s}$ with the initialization $x_{s-1}$ and initial line search value $L_{s-1}\coloneqq L_{{s-1},{N_{s-1}}}.$ $\mathcal{A}_{s}$ will choose $N_{s}$ as $k,$ and output the $k$-th iterate as $x_{s},$ its line search value $L_{s,k}$ when $k\geq 8\left[\tfrac{L_{{s},{k}}(p+\nu)}{4\sigma_{s}}\right]^{\frac{1}{p+\nu}}+1.$ Continue the same number of surplus iterations to record $L_{s,{2N_{s}}}.$
6:   If $\sigma_{s}\geq\tfrac{(L_{s,{2N_{s}}})^{(p+\nu)^{2}}}{(L_{s,{N_{s}}})^{(p+\nu-1)(p+\nu+1)}},$ then terminate with $\widehat{x}=\operatorname*{arg\,min\,}\limits_{N_{s}<k\leq 2N_{s}}\{\|\nabla f(z_{s}^{k})\|\}.$
7: output $\widehat{x}$
Denote $(s,k)$ as the pair referring to the $k$-th iteration of epoch $s,$ and $\mathcal{A}_{s}\coloneqq\mathcal{A}(f,\{\sigma_{i}\}_{i\leq s},\{x_{i-1}\}_{i\leq s}).$ We make the following Assumption on the subroutine $\mathcal{A}_{s}$ used to solve (33).
@@
The subroutine $\mathcal{A}_{s}$ has the following performance guarantee: for any $k>1,$ the following bound holds
@@
where $x_{s}^{k}$ is the computed approximate solution and $L_{s,k}$ is an local estimate of Hölder constant of $f$ at $(s,k)$ such that
@@
where $c_{\mathcal{A}}$ is a universal constant depending on the subroutine $\mathcal{A}_{s},$ and $\theta$ is a user-defined subproblem inexactness parameter used in $\mathcal{A}_{s}.$ For each $s$-th epoch, there exists sequences $\{{z}_{s}^{k}\},$ such that $f({z}_{s}^{k})\leq f(x_{s}^{k})$ and $\nabla f_{s}({z}_{s}^{k}),N_{s}\leq k\leq 2N_{s},$ such that
@@
where $L_{s,2N_{s}}$ is the local estimate of Hölder constant at $(s,2N_{s})$ such that
@@
Notice that 3 relaxes 1 in terms of two aspects. First, it does not require the knowlwdge of $L_{p,\nu}(f),$ the $\nu$-Hölder continuity constant of the $p$-th order derivatives. Instead, it only requires a local estimate $L_{s,k}.$ Second, it allows each iteration of the subroutine $\mathcal{A}_{s}$ to be solved inexactly; this inexactness can be characterized by a parameter $\theta,$ where smaller $\theta$ means a more accurate solution for each iteration. This assumption can be easily satisfied by many algorithms, for example, line search based accelerated tensor methods [14, 15, 16].
@@
Suppose the subroutine satisfies 3. Given $x_{0},L_{0}$ and $\sigma_{1},$ and suppose $\sigma_{1}\leq c_{\mathcal{A}}.$ Furthermore, choose the regularization parameter $\sigma_{s}=2^{p+\nu-1}\sigma_{s-1},$ for all $s\geq 2,$ and the epoch length $N_{s}$ as
(38) $N_{s}\geq 8\left[\tfrac{L_{s,N_{s}}(p+\nu)}{4\sigma_{s}}\right]^{\frac{1}{p+\nu}}+1,\forall s\geq 1.$
@@
after $\mathcal{O}\left(\left[\tfrac{\max\{{pL_{p,\nu}}(f),L_{0},\theta\}}{\sigma_{1}}\right]^{\frac{1}{p+\nu}}\right)$ calls to the $p$-th order oracle.
@@
Similar to the proof of Proposition 2.2, by using the optimality condition of (33), for all $s\geq 1,$ $s\in[S],$ there holds
@@
Combining it with (34) in 3, we have
@@
Combining it with the subroutine sample size condition (38), we have
@@
Similarly, notice that in each epoch $s$, there holds
@@
Combining it with the subroutine sample size condition (38), we have $\|{z_{s}^{k}}-x_{s}^{*}\|\leq\tfrac{1}{4}\|{x_{s-1}}-x_{s-1}^{*}\|.$ Therefore, the conditions of Lemma 3.2 holds, thus, we have
@@
where $x^{*}$ is any solution in the optimal solution set $X^{*}.$ In (i), we substitute the choice for $\sigma_{s}=2^{p+\nu-1}\sigma_{s-1}.$ In (ii), we used $p+\nu\geq 2,$ and $S\geq 1.$ It remains to bound Term I. Substituting the convergence rate of the slow gradient norm from 3 and (41), we have
@@
where in (iii), we used the termination criterion for (4). In (iv), we substitute the linear convergence of $\|x_{s}-x_{s}^{*}\|$ from (43). Substituting the choice of $\sigma_{S}=2^{(p+\nu-1)(S-1)}\sigma_{1},$ we have
@@
Furthermore, by the termination criterion (4), we have
@@
Observe that the RHS is positive due to the condition $\sigma_{1}\leq c_{\mathcal{A}}.$ Note that such a condition is easy to satisfy as one is interested in smaller precision $\sigma_{1},$ and $c_{\mathcal{A}}$ is a constant related to the subroutine $\mathcal{A}.$ The total calls to oracle reads
@@
The expression (39) sheds light on the impact of the initial regularization parameter $\sigma_{1}$ on the convergence. Specifically, the following comments are in order.
@@
    Matching the convergence rate of the function gap residual: If we choose $\sigma_{1}={\varepsilon}/{\{3[9\cdot\text{dist}(x_{0},X^{*})]^{p+\nu-1}\}},$ then Theorem 4.1 shows that Algorithm 4 can compute an approximate solution $\widehat{x}$ s.t. $\|\nabla f(\widehat{x})\|\leq\varepsilon$ after
    (46) $\mathcal{O}\left(\left[\tfrac{\max\{{pL_{p,\nu}}(f),L_{0},\theta\}[C_{p}\text{dist}(x_{0},X^{*})^{p+\nu-1}]}{\varepsilon}\right]^{\frac{1}{p+\nu}}\right)$
    calls to the $p$-th order oracle. This convergence result for the inexact and adaptive gradient minimization is new, and it matched its faster convergence rate for function value residual used in the subroutine $\mathcal{A}$ (cf. 3), see for example, [14, 15, 18, 16]. It improves over the current convergence result for adaptive and inexact gradient minimization as (36) shown in [16].

### v2

@@ 3.1 Extension from accelerated cubic Newton to high-order tensor methods
For the $S$-th epoch, for $N_{S}\leq k\leq 2N_{S},$ there exists $v_{S}^{k}\in\partial f_{S}({x}_{S}^{k}),$ such that
@@
Furthermore, in the last epoch $S,$ for all $N_{S}\leq k\leq 2N_{S},$ we have $\|x_{S}^{k}-x_{S}^{*}\|\leq{\|x_{S-1}-x_{S-1}^{*}\|}/{4}.$
@@
Combining this with (19) in 1 and (17) in Lemma 3.2, we have
@@
Similarly, notice that in the last epoch, for all $N_{S}\leq k\leq 2N_{S},$ we have
@@
Denote $p$-th-order oracle at point $x$ as $\{f(x),\nabla f(x),\nabla^{2}f(x),\dots,\nabla^{p}f(x)\},p\geq 1.$ The following theorem states the main convergence results of the high-order AR framework with subroutine $\mathcal{A}$ applied for convex composite problem (3).
@@
(26) $\mathcal{O}\left(\tfrac{L_{p,\nu}(g)^{\frac{1}{p+\nu}}D^{\frac{p+\nu-1}{p+\nu}}}{\varepsilon^{\frac{1}{p+\nu}}}\right)$
calls to the $p$-th-order oracle.
@@
By Theorem 3.6, we can accelerate the convergence rate of the gradient norm by employing various tensor-based methods as subroutines. Examples include the accelerated cubic Newton method [nesterov2008accelerating], accelerated tensor methods [nesterov2021implementable], the bi-level unconstrained minimization approach [nesterov2023inexact], and accelerated proximal-point methods with lower-order solvers [nesterov2021superfast, nesterov2023inexact]. Notably, in [nesterov2021superfast, nesterov2023inexact], they show that second-order methods can achieve the same convergence rate as third-order methods in terms of the function-value residual. As a result, using such second-order methods as subroutines in the high-order AR framework Algorithm 3 allows the convergence rate in terms of the gradient norm to match that of the function-value residual in [nesterov2021superfast, nesterov2023inexact]. In what follows, we present the Accelerated Regularized Tensor Method (ARTM) [grapiglia2022tensor, grapiglia2020tensor] as an inner subroutine for solving (3) with $\nu\in[0\,1]$, which improves the gradient complexity bound of ARTM in [grapiglia2022tensor, Theorem 5.7] by removing a logarithmic factor. The gradient complexities for the other methods follow by the same reasoning.
@@
Suppose the subroutine $\mathcal{A}$ is ARTM. Then, Algorithm 3 can compute an approximate solution $x_{S}^{k}$ s.t. $\min_{k=N_{S}+1,\dots,2N_{S}}\|\nu_{S}^{k}\|\leq\varepsilon$, where $\nu_{S}^{k}\in\nabla g(x_{S}^{k})+\partial h(x_{S}^{k})$, within $\mathcal{O}(({D^{{p+\nu-1}{}}}/{\varepsilon})^{\frac{1}{p+\nu}})$ calls to the $p$-th-order oracle, where $D$ is defined in (16), and $p\geq 2,$ $\nu\in[0\,\,1].$
@@
[grapiglia2022tensor, Theorem 5.7] shows that 1 holds for ARTM, thus Theorem 3.6 applies.
This result is the tightest known for convex and $p$-times differentiable functions with $\nu$-Hölder continuous $p$-th derivative. Although it does not match the lower complexity bounds for such function class [grapiglia2022tensor, Theorem 6.6], given by $\Omega\left(\left({D^{{p+\nu-1}}}/{\varepsilon}\right)^{\frac{2}{3(p+\nu)-2}}\right).$ The gap arises because no existing methods have yet achieved the optimal rate in terms of the function-value residual for such function class. If such methods are developed, the high-order AR framework can incorporate them as subroutines $\mathcal{A}$ to accelerate the gradient-norm convergence and achieve a matching rate.
@@ 3.2 Fast gradient minimization for convex smooth function
In this subsection, we study AR with optimal $p$-th-order methods as subroutines to handle composite convex functions with Lipschitz continuous $p$-th derivatives. Specifically, there exists $L_{p+1}(g)>0$ s.t.
@@
For the $S$-th epoch, for $N_{S}\leq k\leq 2N_{S},$ there exists $v_{S}^{k}\in\partial f_{S}({x}_{S}^{k}),$ such that
@@
Compared with 1, 2 imposes faster rates for both the function gap and the gradient norm. Note, however, that the gradient-norm rate is still slower than the function-value rate. The above assumption holds for a wide range of optimal algorithms with different $p.$ The subroutines examples include first-order algorithms: the accelerated gradient descent method (AGD) [nesterov2018lectures], second-order algorithms: large stepsize accelerated hybrid proximal extragradient [monteiro2013accelerated], and $p$-th-order algorithms: optimal tensor method (OTM) [bubeck2019near, jiang2021optimal, gasnikov2019optimal] or inexact $p$-th-order proximal point method with tensor step [nesterov2021inexact, nesterov2023inexact].
@@
where $D$ is defined in (16). If the subroutine for solving subproblems satisfies 2. Then, for all $s\in[S],$ Algorithm 3 can compute $x_{s}$ such that $\|x_{s}-x_{s}^{*}\|\leq{\|x_{s-1}-x_{s-1}^{*}\|}/{4}.$ Furthermore, in the last epoch $S,$ for all $N_{S}\leq k\leq 2N_{S},$ we have $\|x_{S}^{k}-x_{S}^{*}\|\leq{\|x_{S-1}-x_{S-1}^{*}\|}/{4}.$
@@
(32) $\mathcal{O}\left(\left[\tfrac{L_{p+1}(g)D^{p}}{\varepsilon}\right]^{\frac{2}{3p+1}}\right)$
calls to the $p$-th-order oracle.
@@
Suppose the subroutine $\mathcal{A}$ is A-NPE [monteiro2013accelerated]. Then, a third-order AR framework Algorithm 3 can compute an approximate solution $x_{S}^{k}$ such that $\min_{k=N_{S}+1,\dots,2N_{S}}\|\nu_{S}^{k}\|\leq\varepsilon$, where $\nu_{S}^{k}\in\nabla g(x_{S}^{k})+\partial h(x_{S}^{k})$, within
@@
[monteiro2013accelerated, Theorem 4.1] shows that 2 holds for A-NPE, thus Theorem 3.11 applies.
We next show a gradient norm minimization using a second-order method for a third-order smooth problem, and show its convergence rate of the gradient norm matches that of the function-value residual in [nesterov2021inexact]. For simplicity, we consider the unconstrained problem (1) in this example.
@@
Suppose the subroutine $\mathcal{A}$ is the second order implementation of the inexact third-order proximal point method [nesterov2021inexact]. Then, a fourth-order AR Algorithm 3 can compute an approximate solution $x_{S}^{k}$ such that
@@
within $\mathcal{O}([{L_{4}(f)D^{3}}/{\varepsilon}]^{\frac{1}{5}})$ iterations, where $D$ is defined in (16), at each iteration, it calls the second-order once and the first-order oracle $\mathcal{O}(\log{\varepsilon^{-1}})$ times at most.
@@
[nesterov2021inexact, Theorem 3.8, 5.5] show that 2 holds with $p=3$, thus Theorem 3.11 applies.
@@ 4 Parameter-free algorithm for convex problems
All convergence results in the Section 2 and Section 3 are based on the exact solutions of the auxiliary optimization problem at each iteration of the high-order method, and assume access to certain key problem parameters, including the Lipschitz constant of the $\nu$-Hölder continuous $p$-th derivative $L_{p,\nu}(f)$, and an upper bound $D$ on the initial distance to the solution set, such that $\operatorname{dist}(x_{0},X^{*})\leq D$. However, in general, the auxiliary subproblem at each iteration cannot be solved exactly, and those problem-dependent parameters are not known beforehand. In this section, we investigate inexact and parameter-free implementations of high-order methods for gradient minimization that do not rely on such information. For clarity of exposition, we first focus on the unconstrained setting; analogous parameter-free strategies can be developed for problems with simple constraints or composite structure.
Denote $(s,k)$ as the pair referring to the $k$-th iteration of epoch $s,$ and $\mathcal{A}_{s}\coloneqq\mathcal{A}(f,\{\sigma_{i}\}_{i\leq s},\{x_{i-1}\}_{i\leq s}).$
Algorithm 4 High-order AR without knowledge of $L_{p,\nu}(f)$
1: Initial regularization $\sigma_{0}=0,\sigma_{1}>0;$ initial point $x_{0}\in\mathbb{R}^{n};$ initial line-search value $L_{0}.$
2: for $s=1,2,\dots$ do
@@
4:   Compute an approximate solution $x_{s}$ of $f_{s}(x)$ by running $\mathcal{A}_{s}$ with the initialization $(x_{s-1},L_{s-1}).$ (33) $x_{s}\approx\operatorname*{arg\,min\,}\limits_{x\in\mathbb{R}^{n}}\left\{f_{s}(x)\coloneqq f(x)+\textstyle\sum_{i=1}^{s}\tfrac{\sigma_{i}-\sigma_{i-1}}{p+\nu}\|x-{x}_{i-1}\|^{p+\nu}\right\},$
5:   and compute $L_{s}=L_{{s-1},{N_{s-1}}},$ the line-search value at $(s-1,N_{s-1}).$ Here $p\geq 1,\nu\in[0,1],$ and $p+\nu\geq 2.$ Let $k$ be the smallest integer satisfying $k\geq 8\left[\tfrac{L_{{s},{k}}(p+\nu)}{4\sigma_{s}}\right]^{\frac{1}{p+\nu}}+1.$
6:   Set $N_{s}\coloneqq k$, and output $x_{s}=x_{s,N_{s}}$ with its line-search value $L_{s,N_{s}}.$
7:   Run an additional $N_{s}$ iterations to record $L_{s,{2N_{s}}}.$
8:   If $\sigma_{s}\geq\tfrac{(L_{s,{2N_{s}}})^{(p+\nu)^{2}}}{(L_{s,{N_{s}}})^{(p+\nu-1)(p+\nu+1)}},$ then terminate with $\widehat{x}=\operatorname*{arg\,min\,}\limits_{N_{s}<k\leq 2N_{s}}\{\|\nabla f(z_{s}^{k})\|\}.$
9: return $\widehat{x}$
We make the following assumption on the subroutine $\mathcal{A}_{s}$ used to solve (33).
@@
The subroutine $\mathcal{A}_{s}$ has the following performance guarantee: for any $k>1,$
@@
where $x_{s}^{k}$ is the computed approximate solution and $L_{s,k}$ is a local estimate of Hölder constant of $f$ at $(s,k)$ such that
@@
where $c_{\mathcal{A}}$ is a universal constant depending on the subroutine $\mathcal{A}_{s},$ and $\theta$ is a user-defined subproblem inexactness parameter controlling the inner accuracy in $\mathcal{A}_{s}.$ For each $s$-th epoch, for $N_{s}\leq k\leq 2N_{s},$ there exists sequences $\{{z}_{s}^{k}\}$ with $f({z}_{s}^{k})\leq f(x_{s}^{k})$ and $\nabla f_{s}({z}_{s}^{k}),$ such that
@@
where $L_{s,2N_{s}}$ is a local estimate of Hölder constant at $(s,2N_{s})$ such that
@@
Notice that 3 relaxes 1 in two aspects. First, it does not require knowledge of $L_{p,\nu}(f);$ instead, it uses local estimate $L_{s,k}.$ Second, it allows inexact inner solution, quantified by $\theta,$ where smaller $\theta$ means a more accurate solution for each iteration. This assumption can be easily satisfied by many adaptive algorithms, for example accelerated tensor methods [grapiglia2020tensor, grapiglia2022tensor, grapiglia2023adaptive, He2025HAR].
@@
Suppose the subroutine satisfies 3. Given $x_{0},L_{0}$ and $\sigma_{1},$ and suppose $\sigma_{1}\leq c_{\mathcal{A}}.$ Furthermore, choose the regularization parameter $\sigma_{s}=2^{p+\nu-1}\sigma_{s-1},$ for all $s\geq 2.$ Suppose the epoch length $N_{s}$ as
(38) $N_{s}\geq 8\left[\tfrac{L_{s,N_{s}}(p+\nu)}{4\sigma_{s}}\right]^{\frac{1}{p+\nu}}+1,\quad\forall s\geq 1.$
@@
after $\mathcal{O}\left(\left[\tfrac{\max\{{pL_{p,\nu}}(f),L_{0},\theta\}}{\sigma_{1}}\right]^{\frac{1}{p+\nu}}\right)$ calls to the $p$-th-order oracle.
@@
Similar to the proof of Proposition 2.2, by using the optimality condition of (33), for all $s\geq 1,$ $s\in[S],$ we have
@@
Combining this with (34) in 3, we have
@@
Combining this with the subroutine sample size condition (38), we have
@@
Similarly, notice that in each epoch $s$, we have
@@
Combining this with the subroutine sample size condition (38), we have $\|{z_{s}^{k}}-x_{s}^{*}\|\leq\tfrac{1}{4}\|{x_{s-1}}-x_{s-1}^{*}\|.$ Therefore, the conditions of Lemma 3.2 holds, thus, we have
@@
where $x^{*}$ is any solution in the optimal solution set $X^{*}.$ In (i), we substitute the choice for $\sigma_{s}=2^{p+\nu-1}\sigma_{s-1}.$ In (ii), we used $p+\nu\geq 2,$ and $S\geq 1.$ It remains to bound Term I. Substituting the slow gradient norm convergence rate from 3, we have
@@
where in (iii), we used the termination rule in Algorithm 4. In (iv), we substitute the linear convergence of $\|x_{s}-x_{s}^{*}\|$ from (43). Substituting the choice of $\sigma_{S}=2^{(p+\nu-1)(S-1)}\sigma_{1},$ we have
@@
Furthermore, by the termination rule in Algorithm 4, we have
@@
The RHS is positive as $\sigma_{1}\leq c_{\mathcal{A}}.$ This is mild as $\sigma_{1}$ serves as the accuracy precision (small) and $c_{\mathcal{A}}$ is a subroutine-dependent constant. The total oracle calls read
@@
The expression (39) clarifies the dependence of the convergence rate on the initial regularization parameter $\sigma_{1}.$ Specifically, the following comments are in order.
@@
    Matching the convergence rate of the function-value: If we choose $\sigma_{1}={\varepsilon}/{\{3[9\cdot\operatorname{dist}(x_{0},X^{*})]^{p+\nu-1}\}},$ then by Theorem 4.1, Algorithm 4 can compute an approximate solution $\widehat{x}$ such that $\|\nabla f(\widehat{x})\|\leq\varepsilon$ after
    (46) $\mathcal{O}\left(\left[\tfrac{\max\{{pL_{p,\nu}}(f),L_{0},\theta\}\operatorname{dist}(x_{0},X^{*})^{p+\nu-1}}{\varepsilon}\right]^{\frac{1}{p+\nu}}\right)$
    calls to the $p$-th-order oracle. This convergence result for inexact and adaptive gradient minimization is new and matches the faster function-value rate in the subroutine $\mathcal{A}$ (cf. 3), see for example, [grapiglia2020tensor, grapiglia2022tensor, jiang2020unified, grapiglia2023adaptive]. It improves over the current convergence result for adaptive and inexact gradient minimization as (36) shown in [grapiglia2023adaptive].